1. Below are the data of Milk and Honey in an economy
YEAR PRICE OF Milk QUANTITY OF Milk PRICE OF Honey QUANTITY OF Honey
2017 (Base year) $1 100 $2 50
2018 $1 200 $2 100
2019 $2 200 $4 100

a) Compute nominal GDP, real GDP and GDP deflator for each year.
b) Compute the percentage change in nominal GDP, real GDP and GDP deflator in 2018 and 2019 from the preceding year.
1
Expert's answer
2020-04-13T09:08:02-0400

"a) nominal GDP" 2017"=1\u00d7100+2\u00d750=" $"200;"

"nominal GDP" 2018"=" "1\u00d7200+2\u00d7100=" $"400;"

"nominal GDP" 2019 "=2\u00d7400+4\u00d7100=" $ "800;"

"real GDP" 2017"=1\u00d7100+2\u00d750=" $"200;"

"real GDP" 2018"=1\u00d7200+2\u00d7100=" $"400;"

"realGDP" 2019"=1\u00d7200+2\u00d7100=" $"400;"

GDP deflator2017"=200\/200=1;"

GDP deflator2018"=400\/400=1;"

GDP deflator2019"=800\/400=2;"

"b)" change in nominal GDP2018"=\frac {400-200}{200}×100" "=" %"100;"

change in nominal GDP2019 "=\frac {800-400}{400}×100=" %"100;"

change in real GDP2018 "=\frac{400-200}{200}×100=" %100;

change in real GDP2019"=\frac {400-400}{400}×100=" %0;

change in GDP deflator2018"=\frac {1-1}{1}×100=" %"0;"

change in GDP deflator2019"=\frac{2-1}{1}×100=" %"100;"